OpenStudy (niahloves):

Which property justifies this statement? If x + y = z and z = x + 4, then x + y = x + 4. A. symmetric B. associative C. distributive D. transitive

OpenStudy (niahloves):

@LynFran

OpenStudy (thomas5267):

D. https://en.wikipedia.org/wiki/Transitive_relation "In mathematics, a binary relation R over a set X is transitive if whenever an element a is related to an element b, and b is in turn related to an element c, then a is also related to c."

OpenStudy (niahloves):

what's the answer

OpenStudy (thomas5267):

The answer is D. Symmetric property is when aRb if and only if bRa where R is a relation on some numbers. For example, equality is a symmetric relations since a=b if and only if b=a. An example of a relation that is not symmetric is <. If a<b is true then b<a is not true.\ Associative property is related to operations on number. It means that (a+b)+c=a+(b+c) for some binary operation +. For example, ordinary multiplication and addition are associative operations. Distributive property is related to two operations. It means that (a+b)*c=ac+bc for some operation + and *. Ordinary multiplication distributes over addition since a*(b+c)=ab+ac. Transitive property is when aRb and bRc implies aRc for some relation R. Equality is a transitive relation since a=b and b=c implies a=c.